Find the coordinate of the circle of the centre and its radius if its equation is

(a) x² + y² - 12y + 6 = 0

(b) -2x² - 2y² - 18y + 9 = 0


1
Expert's answer
2021-06-14T16:10:32-0400

Answer:-

general equation : x2+y2+2gx+2fy+c=0

(a) x² + y² - 12y + 6 = 0

Centre= (-g,-f)= (0,6)

Radius="\sqrt{g^2+f^2-c}" ="\sqrt{0+36-6}"

=5.4

(b) -2x² - 2y² - 18y + 9 = 0

deciding by -2

We get x2+y2+9y+4.5=0

Centre= (-g,-f)= (0,-4.5)

Radius="\sqrt{g^2+f^2-c}"

="\sqrt{0+(4.5)^2-4.5}"

=3.96